# Gaussian Process Regression as a Sustainable Data-driven Background Estimate Method at the (HL)-LHC

### v2

@@ 1 Introduction
Gaussian Process Regression (GPR), on the other hand, is a non-parametric technique broadly used in machine learning and statistics [26]. A Gaussian Process is a collection of random variables such that any finite combination of them has a joint Gaussian distribution. Its potential usage in HEP experiments is discussed in ref. [27, 28] and several LHC results have applied this method to achieve various goals, such as template smoothing [29] or background estimation [18, 30]. In a GPR, the bin contents $\nu_{1}...\nu_{n}$ are described by a Gaussian PDF:
$p(\nu_{i};\mu_{i},\mathbf{K})=\frac{1}{\sqrt{(2\pi)^{n}|\mathbf{K}|}}\exp\left(-\frac{1}{2}\sum_{i,j}^{n}(\nu_{i}-\mu_{i})K^{-1}_{ij}(\nu_{j}-\mu_{j})\right),$
where $\mu_{i}$ is the bin centre and $\mathbf{K}$ is the covariance matrix, which is parameterised by a kernel function $K(x_{i},x_{j})$. A common choice of kernel is the Radial Basis Function (RBF) kernel:
$K(u_{i},u_{j})=\exp\left(-\frac{\lvert\lvert u_{i}-u_{j}\rvert\rvert^{2}}{2\ell^{2}}\right),$
where $\ell$ is the length scale of the kernel and $u_{i(j)}$ refers to the bin centre of the training data. It determines how closely correlated the neighbouring points are and can be fit to data through maximising the marginal likelihood of the observations. The authors of ref. [27] and ref. [28] explored the application of GPR in di-jet resonance searches. While previous work was concentrated on testing different kernels, in this work, more attention is paid to the hyperparameters, the pre-defined parameters that are not optimised by GPR. It is demonstrated that the $L_{2}$ regularisation, which represents the level of noise in the input dataset, can be introduced as a set of hyperparameters to control the overall performance of GPR. Applying the widely adopted RBF kernel, with minimal hyperparameter tuning, GPR is capable of fitting a complicated spectrum that is challenging for functional fit methods. It is also robust against the increasing luminosity up to the HL-LHC era.
@@ 3 Model Setup
This work utilises the GPR module available in Scikit-learn 1.3.0, without any modifications to the core libraries.
@@ 3.1 Data Pre-processing
The dataset used for resonance searches in the hadronic final states is often very sizeable. In this work, we use a binned dataset, so for each bin centre ($m_{\mathrm{jj}}$) there is a certain number of entries ($N_{\mathrm{events}}$), corresponding to the input data points ($x$) and the ones to be predicted by GPR ($y$), respectively. Its rapidly changing nature and the large mass coverage require a proper pre-processing step. Applying the logarithms of $x$ and $y$ can achieve both fast convergence and good performance. In the case of empty bins where $\log(y_{i})$ is not defined, zero padding is adopted. Figure 2 shows an example pre-processed background dataset and hypothetical Gaussian-shaped signal events. A Gaussian-shaped signal approximates a resonance with a given mass at the mean. The width of the resonance is parametrised by the ratio between the standard deviation and the mean. A benchmark value of 5% is chosen as it is a frequently considered case. The signal is well contained within an approximately 0.5 interval on the $x$-axis, motivating the hyperparameter choice made in Section 3.4.
Figure 2: An example of the pre-processed background-only pseudo-datasets (left) and an illustration of the distributions for Gaussian-shaped signal at the 250, 300, 800 and 1400 GeV mass points, with a 5% width (right).
3.2 Kernel Selection
The kernel choice has a significant impact on the performance. As already mentioned in Section 1, there is a large series of kernels to choose from [26]. The goal of this work is not to find or construct the most optimal kernel, but to demonstrate the generality of a GPR-based method. Therefore, we use the multiplication of the two most common ones, an RBF and a constant kernel: $\mathrm{C}(c)\times\mathrm{RBF}(\ell)$. Previous works [27, 28] have shown this combination is a viable choice for HEP experiments. The only two parameters to be optimised are $c$ and $\ell$, and the bounds for $c$ and $\ell$ are hyperparameters defined a priori. The length scale of the RBF kernel, $\ell$, determines the smoothness of the model, and it has to be large enough in order not to incorporate the signal events. It is because a localised feature like a narrow signal will create bumpy features.
@@ 3.3 Regularisation
In the functional fit method, users have a set of criteria, such as what is listed in Section 4. When a function cannot satisfy this list, usually the initial reaction is to find a better-performing one. Similarly, we can also try to craft a more suitable customised kernel in a GPR-based strategy. However, this problem can be addressed differently, by biasing the minimisation to satisfy the imposed standards.
@@ 4 Background Estimate Validation
A background modelling method should be validated with background-only samples. Well-known goodness-of-fit tests, such as the reduced $\chi^{2}$ test [33], can quantify the modelling performance, where a $\chi^{2}/\mathrm{nDoF}$ close to unity indicates a well-behaved background estimate. In addition, the Kolmogorov–Smirnov (KS) test evaluates how compatible the significance is with a normal distribution [34], as it is expected when the background estimate is unbiased. A KS $p$-value smaller than the threshold rejects the hypothesis that the significance is a normal distribution. An ensemble test using pseudo-experiments tells us the fraction of trials that gives acceptable reduced $\chi^{2}$ or KS test result. Though these two classic tests reveal the quality of the overall performance, they cannot detect local biases efficiently. As the search strategy is designed to find narrow peaks, such mis-modelling is likely to induce false-positive errors.
@@ 5.1 Sensitivity Evaluation
$S=Ae^{-\frac{1}{2}(\mu_{i}-\mu_{j})^{2}/{l_{s}}^{2}}e^{-\frac{1}{2}((\mu_{i}-m)^{2}+(\mu_{j}-m)^{2})/t^{2}}$
where $A$ is a constant, $l_{s}$ refers to the length scale, $m$ specifies the centre of the signal mass, and $t$ acts as the width of the signal [27]. Therefore, the full GPR model becomes $S+\mathrm{C}(c)\times\mathrm{RBF}(\ell)$. The pseudo-experiments done in Section 4 using background-only spectra are used to determine the bounds of the length scale of the background kernel. Most of those pseudo-experiments report a best-fitted length scale between 0.6 and 0.7, so $l_{0}$ ($l_{1}$) is set to be 0.6(0.7). Similarly, the hyperparameters associated with the signal kernel are determined by fitting the signal templates. While both $A$ and $l_{s}$ can float freely, $m$ and $t$ are only allowed to change within a reasonable range of the fitted values. Given the logarithmic transformation done to the dataset in Section 3.1, the signal in the original space is approximated by $e^{s_{i}+b_{i}}-e^{b_{i}}$, where $s_{i}$ and $b_{i}$ are the signal and background components in the $i$-th bin, predicted by GPR. Figure 9 shows the signal extraction results for the 800 GeV Gaussian-shaped signal with a 5% width. The number of extracted signal events increases as the amount of injected signal events, roughly following a linear response. However, the extracted signal strength is systematically higher and has a wide spread. The performance of signal extraction becomes very unstable for signal points close to the $m_{\mathrm{jj}}$ boundaries. The data pre-processing, kernel selection and hyperparameters all affect the signal extraction performance, which can be further optimised. We use this example to demonstrate there exists a viable strategy.

### v3

@@ 1 Introduction
Gaussian Process Regression (GPR), on the other hand, is a non-parametric technique broadly used in machine learning and statistics [26]. A Gaussian Process is a collection of random variables such that any finite combination of them has a joint Gaussian distribution. Its potential usage in HEP experiments is discussed in ref. [27, 28] and several LHC results have applied this method to achieve various goals, such as template smoothing [29] or background estimation [18, 30]. In a GPR, the bin contents $y_{1}...y_{n}$ are described by a Gaussian PDF:
$p(y_{i};\mu_{i},\mathbf{K})=\frac{1}{\sqrt{(2\pi)^{n}|\mathbf{K}|}}\exp\left(-\frac{1}{2}\sum_{i,j}^{n}(y_{i}-\mu_{i})K^{-1}_{ij}(y_{j}-\mu_{j})\right),$
where $\mu_{i}$ is the prior mean of $y_{i}$, and $\mathbf{K}$ is the covariance matrix, parameterised by a kernel function $K(x_{i},x_{j})$. A common choice of kernel is the Radial Basis Function (RBF) kernel:
$K(x_{i},x_{j})=\exp\left(-\frac{\lvert\lvert x_{i}-x_{j}\rvert\rvert^{2}}{2\ell^{2}}\right),$
where $\ell$ is the length scale of the kernel and $x_{i(j)}$ refers to the bin centre of the training data. It determines how closely correlated the neighbouring points are and can be fit to data through maximising the marginal likelihood of the observations. The fitted model gives us the posterior distribution of $y_{i}$, which can be used to calculate the mean and the standard deviation of the estimate. The authors of ref. [27] and ref. [28] explored the application of GPR in di-jet resonance searches. While previous work was concentrated on testing different kernels, in this work, more attention is paid to the hyperparameters, the pre-defined parameters that are not optimised by GPR. It is demonstrated that the $L_{2}$ regularisation, which represents the level of noise in the input dataset, can be introduced as a set of hyperparameters to control the overall performance of GPR. Applying the widely adopted RBF kernel, with minimal hyperparameter tuning, GPR is capable of fitting a complicated spectrum that is challenging for functional fit methods. It is also robust against the increasing luminosity up to the HL-LHC era.
@@ 3 Model Setup
This work utilises the GPR module available in Scikit-learn 1.5.1, without any modifications to the core libraries.
@@ 3.1 Data Pre-processing
The dataset used for resonance searches in the hadronic final states is often very sizeable. In this work, we use a binned dataset, so for each bin centre ($m_{\mathrm{jj}}$) there is a certain number of entries ($N_{\mathrm{events}}$), corresponding to the input data points ($x$) and the ones to be predicted by GPR ($y$), respectively. Its rapidly changing nature and the large mass coverage require a dedicated pre-processing step. Applying the logarithms of $x$ and $y$ can achieve both fast convergence and good performance. In the case of empty bins where $\log(y_{i})$ is not defined, zero padding is adopted. Figure 2 shows an example pre-processed background dataset and hypothetical Gaussian-shaped signal events. A Gaussian-shaped signal approximates a resonance with a given mass at the mean. The width of the resonance is parametrised by the ratio between the standard deviation and the mean. A benchmark value of 5% is chosen as it is a frequently considered scenario in resonance searches. The signal is well contained within an approximately 0.5 interval on the $x$-axis, motivating the hyperparameter choice made in Section 3.4.
Figure 2: An example of the pre-processed background-only pseudo-datasets (left) and an illustration of the distributions for Gaussian-shaped signal at the 250, 300, 800 and 1400 GeV mass points, with a 5% width (right). Each signal point is normalised to 10K events.
3.2 GPR Model Setup
The kernel choice has a significant impact on the performance. As already mentioned in Section 1, there is a large series of kernels to choose from [26]. The goal of this work is not to find or construct the most optimal kernel, but to demonstrate the generality of a GPR-based method. Therefore, we use the multiplication of the two most common ones, an RBF and a constant kernel: $\mathrm{C}(c)\times\mathrm{RBF}(\ell)$. Previous works [27, 28] have shown this combination is a viable choice for HEP experiments. The only two parameters to be optimised are $c$ and $\ell$, and the bounds for $c$ and $\ell$ are hyperparameters defined a priori. The length scale of the RBF kernel, $\ell$, determines the smoothness of the model, and it has to be large enough in order not to incorporate the signal events. It is because a localised feature like a narrow signal will create bumpy features. The prior mean of $y_{i}$ has a diminishing impact on the posterior distribution, and is set to zero.
@@ 3.3 Regularisation
For the functional fit method, users have a set of criteria, such as what is listed in Section 4. When a function cannot satisfy those criteria, usually the initial reaction is to find a better-performing one. Similarly, we can also try to craft a more suitable customised kernel in a GPR-based strategy. However, this problem can be addressed differently, by biasing the minimisation to meet the imposed standards.
@@ 4 Background Estimate Validation
A background modelling method should be validated with background-only samples. Well-known goodness-of-fit tests, such as the reduced $\chi^{2}$ test [33], can quantify the modelling performance, where a $\chi^{2}/\mathrm{nDoF}$ close to unity indicates a well-behaved background estimate. In addition, the Kolmogorov–Smirnov (KS) test evaluates how compatible the significance is with a normal distribution [34], as it is expected when the background estimate is unbiased. A KS $p$-value smaller than the threshold rejects the hypothesis that the significance is a normal distribution. An ensemble test using pseudo-experiments tells us the fraction of trials that gives an acceptable reduced $\chi^{2}$ or KS test result. Though these two classic tests reveal the quality of the overall quality, they cannot detect local biases efficiently. As the search strategy is designed to find narrow peaks, such mis-modelling is likely to induce false-positive errors.
@@ 5.1 Sensitivity Evaluation
$S=Ae^{-\frac{1}{2}(x_{i}-x_{j})^{2}/{l_{s}}^{2}}e^{-\frac{1}{2}((x_{i}-m)^{2}+(x_{j}-m)^{2})/t^{2}}$
where $A$ is a constant, $l_{s}$ refers to the length scale, $m$ specifies the centre of the signal mass, and $t$ acts as the width of the signal [27]. Therefore, the full GPR model becomes $S+\mathrm{C}(c)\times\mathrm{RBF}(\ell)$. The pseudo-experiments done in Section 4 using background-only spectra are used to determine the bounds of the length scale of the background kernel. Most of those pseudo-experiments report a best-fitted length scale between 0.6 and 0.7, so $l_{0}$ ($l_{1}$) is set to be 0.6(0.7). Similarly, the hyperparameters associated with the signal kernel are determined by fitting the signal templates directly. While both $A$ and $l_{s}$ can float freely, $m$ and $t$ are only allowed to change within a reasonable range of the fitted values. Given the logarithmic transformation done to the dataset in Section 3.1, the signal in the original space is approximated by $e^{s_{i}+b_{i}}-e^{b_{i}}$, where $s_{i}$ and $b_{i}$ are the signal and background components in the $i$-th bin, predicted by GPR. Figure 9 shows the signal extraction results for the 800 GeV Gaussian-shaped signal with a 5% width. The number of extracted signal events increases as the amount of injected signal events, roughly following a linear response. However, the extracted signal strength is systematically higher and has a wide spread. The performance of signal extraction becomes very unstable for signal points close to the $m_{\mathrm{jj}}$ boundaries. The data pre-processing, kernel selection and hyperparameters all affect the signal extraction performance, which can be further optimised.